Power Temporaries Inc. has earnings of $9,000,000 with 2,100,000 shares outstanding before a public distribution. Eleven hundred thousand shares will be included in the sale, of which 700,000 are new corporate shares and 400,000 are shares currently owned by Julie Lipner, the founder and CEO. The 400,000 shares that Julie is selling are referred to as a secondary offering and all proceeds will go to her. (Round the final answers to 2 decimal places.)

a. What were the corporation’s earnings per share before the offering? Earnings per share $

b. What are the corporation’s earnings per share expected to be after the offering? Earnings per share $
Answer #1

a)Earning per share =[Net income -preferred dividend if any]/number of common shares outstanding

                   = [9,000,000 -0 ]/2,100,000

                   = 9,000,000 /2,100,000

                    = $ 4.29 per share

b)Number of shares outstanding after new offerings = 2,100,000 + 700,000 = 2,800,000shares.

Shares traded in secondary market are not considered as offerings by corporate that is only new shares issued by corporate to public or initial public offerings are considered while calculating number of shares outstanding.

400000 shares traded or sold are already from existing shares and not from new offerings /issuance and thus are not considered while calculating number of shares outstanding

Earning per share = 9,000,000 / 2,800,000

                     = $ 3.21 per share